OpenStudy (anonymous):

prove by induction that for all integral values of n 2^(n+2) + 3^(2n+1) is divisible by 7

OpenStudy (anonymous):

in equation form \[2^{n+2} + 3^{2n+1} \] is divisible by 7 for integral n

jhonyy9 (jhonyy9):

so this is very easy first calcule for n=0 and result that is true so for n=1 make calcule and result that is true after this suppose that for n=k is true and after this you need to prove that this will be true for n=k+1 too

jhonyy9 (jhonyy9):

those are all by proving by induction method !!! good luck ! bye

OpenStudy (anonymous):

yep, i understnd the induction method quite well, but proving the n=k+1 has had me going around in circles for about an hour!

OpenStudy (anonymous):

i can see that its \[2^{k+3} + 3^{2k+3}\]

OpenStudy (anonymous):

its how to factor out the 7

OpenStudy (slaaibak):

proof for n=1: 2^3 + 3^3 = 8 + 27 = 35 Therefore for n=1 it is divisible by 7. Now suppose \[2^{n+2} + 3^{2n + 1}\] is divisible by 7. Prove it for n+1: \[2^{n + 1 + 2} + 3^{2(n+1) + 1}\] \[=2^{n+3} + 3^{2n + 3}\] Write it like this: \[2 * 2^{n+2} + 3^2 * 3^{2n+2}\] Factor out the 2: \[2(2^{n+2} + 3^{2n + 1} ) + 3^2*3^{2n+1} - 2 * 3^{2n+1}\] Factorize: \[2(2^{n+2} + 3^{2n+1}) + 3^{2n+1} * (3^2 - 2)\] \[2(2^{n+2} + 3^{2n+1}) + 3^{2n+1} * 7\] Since: \[2(2^{n+2} 3^{2n+1} )\] is divisible by 7, since we said Suppose: \[ ({} 2^{n+2} + 3^{2n + 1})\] is divisible by 7. now, the second part is also divisible by 7: \[3^{2n+1} * 7\] Because the term is multiplied by 7. Therefore, because both terms are divisible by 7, this is true for n+1 and therefore all n >=1 This completes the proof.

OpenStudy (slaaibak):

It should be \[2(2^{n+2} + 3^{2n + 1})\] and not \[2(2^{n+2}3^{2n+1})\] I made an error in the one line

OpenStudy (anonymous):

briliant- thanks !

OpenStudy (slaaibak):

Shot